OpenStudy (anonymous):

the derivate of h(x)=4/7x is 4/7?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

ok i got it now

OpenStudy (anonymous):

jimmy can you help me with this find the slope of the tangent line of f(x)= 2x^2-x+3

OpenStudy (anonymous):

best written as (6/7) x though jinnie - so its not confused with 4 divided by 7x

OpenStudy (anonymous):

6/7?

OpenStudy (anonymous):

yea ok find derivative of f(x) f'(x) = 4x - 1 - thats it

OpenStudy (anonymous):

so for example the slope of tangent at x = 1 = 4(1) - 1 = 3

OpenStudy (anonymous):

sorry i meant (4/7)

OpenStudy (anonymous):

so i must find the derivative first which is 2x-1+0

OpenStudy (anonymous):

of 2x^2-x+3

OpenStudy (anonymous):

yeah - but its 4x - 1 The derivative at a point on a curve is slope of the tangent at that point

OpenStudy (anonymous):

oh yea you have to multiply ok i got it

OpenStudy (anonymous):

so what is the equation of the tangent line at x=1? is it the same as the slope of the tangent line at the point x=1?

OpenStudy (anonymous):

no use the formula y-y1 = m(x-x1) to get the equation at x=1 , y = 2(1)^2 - 1 + 3 = 4 so we need the equation at the point (1,4) with a slope of 3 y - 4 = 3 (x - 1) is equation of the tangent

OpenStudy (anonymous):

can i continue with that or is that where it stops

OpenStudy (anonymous):

you can simplify that y - 4 = 3x - 3 y = 3x + 1

OpenStudy (anonymous):

so the way to find the equation of the tangent at a point on a curve is to 1. differentiate the equation of the curve to find slope of the curve 2 substitute for slope, and x and y values of the point into the general form y-y1 = m(x-x1)

OpenStudy (anonymous):

alright thanks

OpenStudy (anonymous):

how did you get the slope of 3 btw

OpenStudy (anonymous):

yw - hope this helps

OpenStudy (anonymous):

i found the slope in terms of x by differentiating f(x) = 2x^2 - x +3 f'(x) = 4x - 1 - this the slope in terms of x , at point x=x so at x = 1, slope is found by plugging x=1 into 4x-1: slope = 4(1) - 1 = 3

OpenStudy (anonymous):

ok thanks